5. (x + 2y − 1)dx − (2x + y − 5)dy = 0


1
Expert's answer
2021-05-11T15:05:05-0400

The give differential equation is-

(x + 2y − 1)dx − (2x + y − 5)dy = 0


first, re-arrange the problem to be-


"x+2y-1+(2x+y-5)\dfrac{dy}{dx}=0"


let "M=x+2y-1" and "N=2x+y-5" and define "Fx=M" and "Fy=N"


Differentiate M w.r.t y-

"M_y=2"


Differentiate N w.r.t. x-

"N_x=2"


As "M_y=N_x"

thus the problem is an EXACT ODE


fourth, find integral of M dx holding y constant "= \dfrac{x^2}{2}+2xy-x+h(y) = F"


find "F_y= 2x+h'= N = 2x+y-5 ..." ' solve for "h'=y-5" and "h=\dfrac{y^2}{2}-5y"


Complete solution is -


"\dfrac{x^2}{2}+2xy-x+\dfrac{y^2}{2}-5y=c"